OpenStudy (anonymous):

Please help why I am getting another answer. I would like to know if there is an easy way to solve this or at least with th calculator......Jijiji!! ∫[1,4,(t^(7)-t^(3))/(t^(5))dx] Thanks!!

zepdrix (zepdrix):

Is 1 the lower limit? Try to get in the habit of writing the LOWER limit first ^^ 1,4 Cheater way using a calculator? Hmmm I'm not sure :) lol This one is pretty straight forward though. You just have to simplify your fraction, then you can apply the power rule for integration.

OpenStudy (anonymous):

I just found the mistake..

OpenStudy (anonymous):

it supposed to be this one..∫[1,4,(t^(7)-t^(3))/(t^(5)),]

zepdrix (zepdrix):

\[\large \int\limits_1^4 \frac{t^7-t^3}{t^5}dt \quad = \quad \int\limits_1^4 \frac{t^7}{t^5}-\frac{t^3}{t^5}dt \quad = \quad \int\limits_1^4 t^{7-5}-t^{3-5}dt\]

zepdrix (zepdrix):

Oh did you figure it out already? :D

OpenStudy (anonymous):

Nop!! I thought so.. :(

OpenStudy (anonymous):

I still don't get it!!

zepdrix (zepdrix):

Ahh sorry I musta missed the blinky message up in the corner D:::

zepdrix (zepdrix):

Understand what I did with the powers on t??

OpenStudy (anonymous):

No, I believe you substracted them.

zepdrix (zepdrix):

Yah when you divide terms with the same base, you SUBTRACT the exponents c: You reposted this thread it looks like though.. so I assume you figured this out maybe? :D

OpenStudy (anonymous):

Yes I did, but I have a concern, I have 1/3(4^3 -1) +(1/4-1) where they coming from.. How its simplify?

OpenStudy (anonymous):

At least the fractions.. 1/3 and 1/4..

zepdrix (zepdrix):

\[\large \int\limits\limits_1^4 t^{7-5}-t^{3-5}dt \quad = \quad \int\limits_1^4 t^2-t^{-2}dt\]Integrating gives us,\[\large \frac{1}{3}t^3+t^{-1} \quad |_1^4 \quad \rightarrow \qquad \frac{1}{3}t^3+\frac{1}{t}\quad|_1^4\]

zepdrix (zepdrix):

\[\large \left(\frac{64}{3}+\frac{1}{4}\right)-\left(\frac{1}{3}+1\right)\]

zepdrix (zepdrix):

I think you get something like that, hopefully I didn't mess anything up...

OpenStudy (anonymous):

Thanks!!